Find the length and width of a rectangle whose perimeter is 36 feet and whose area is 80 square feet.
Answer by stanbon(75887)

Find the length and width of a rectangle whose perimeter
is 36 feet and whose area is 80 square feet.
------
Perimeter Eq: 36 = 2(L+W)
Area Eq; LW = 80
--------
Modify:
L + W = 18
----
L(18-L) = 80
-L^2+18L-80 = 0
----
Use The Quadratic Formula to solve for "L":
L = [-18 +- sqrt(18^2-4*-1*-80)]/(-2)
---
L = [-18 +- sqrt(4)]/-2
---
Positive solution:
L = [-18-2]/-2 = 10 (length)
---
Since L+W = 18, W = 8 (width)
==================================
Cheers,
Stan H.
=========